# Joyce structures from quadratic differentials on the sphere

### v1

@@ 1. Introduction
Joyce structures were proposed by Bridgeland in [8] as a way of geometrising the study of Donaldson-Thomas invariants. They are expected to exist on the complex manifold $M$ of stability conditions of a $CY_{3}$ triangulated category satisfying certain assumptions. The argument for their existence involves solving Riemann-Hilbert problems with infinite dimensional gauge group, defined by the variation of the Donaldson-Thomas invariants and as far as the author understands, solutions to these problems are not yet fully understood. A geometric axiomatisation of Joyce structures without reference to their origin in Donaldson-Thomas theory is provided in [12] in terms of a complex hyper-Kähler metric on the total space $TM$ satisfying certain symmetries compatible with an integral structure on $M$.
@@
The symplectic geometry and Hamiltonian structure of isomonodromic deformations has long been studied [3] and is still of active interest [2]. The infinitesimal generators may be obtained as the kernel of a fundamental 2-form and formulae in a general setting are provided in [27]. In this work, within the setting of second order linear ODEs in one dependent variable, we take a (at least superficially) different approach and construct a fundamental $2$-form in terms of the intersection pairings of a family of algebraic curves.
We will calculate isomonodromic deformations of a large class of ODEs in complex variable $x$ of the form
$\frac{d^{2}Y}{dx^{2}}=Q(x)Y,$ (1.1)
where the potential takes the form
$Q(x)=\frac{Q_{0}(x)}{\hbar^{2}}+\frac{Q_{1}(x)}{\hbar}+Q_{2}(x)$ (1.2)
where $Q_{0}(x),Q_{1}(x),Q_{2}(x)$ are rational functions depending on a point $\xi\in X$ where $X$ is a complex manifold, and $\hbar\in\mathbb{C}^{*}$ is the spectral parameter.
@@ 2. Complex hyper-Kähler metrics and their twistor theory
Let $X$ be a complex manifold of dimension $4n$. A complex hyper-Kähler structure is a section $g$ of $\odot^{2}T^{*}X$, where $TX$ denotes the holomorphic tangent bundle of $X$, defining a non-degenerate bilinear form on each fibre of $TX$, with holomorphic endomorphisms $I,J,K$ of $TX$ satisfying
@@
$I^{*}g$ $=J^{*}g=K^{*}g=g$
$\nabla I$ $=\nabla J=\nabla K=0$
@@
Crucial to constructing our metrics will be families of rank-$2n$ subbundles of $TX$ depending on a spectral parameter $\hbar=u^{1}/u^{0}\in\mathbb{C}^{*}$ where $u^{0}:u^{1}]\in\mathbb{CP}^{1}$
@@
$L(\hbar)=\operatorname{span}\Big{\{}L_{a}\coloneqq U_{a}+\frac{V_{a}}{\hbar}\Big{\}}_{a=1}^{2n}$ (2.2)
@@
$g=\sum_{a=1}^{2n}\sum_{b=1}^{2n}\omega_{ab}(U^{a}\otimes V^{b}+V^{a}\otimes U^{b})$ (2.4)
@@
$L(\hbar)$ Frobenius integrable for all $\hbar$ is equivalent to integrability of the $\pm i$-eigenspaces of the complex structures $I,J,K$. We may think of $L$ as a subbundle of $T(X\times\mathbb{CP}^{1})$. Then, integrability implies we may construct the space of leaves of this distribution.
@@
$\wedge^{k}_{p}Z\coloneqq\wedge^{k}({\operatorname{ker}dp})^{*},$ (2.6)
denote the bundle of $k$-forms relative to $p$. A relative (to $p$) $\mathcal{O}(2)$-valued 2-form is a section of the bundle $\wedge^{2}_{p}Z\otimes p^{*}\mathcal{O}(2)$.
We have the relative exterior derivative $d_{p}:\Gamma(\wedge^{2}_{p}Z)\to\Gamma(\wedge^{3}_{p}Z)$ given by
$d_{p}\omega=d\tilde{\omega}|_{\ker p}$ (2.7)
where $\tilde{\omega}$ is any $2$-form that restricts to $\omega$ on $\operatorname{ker}p$. If $d_{p}\omega=0$ we say $\omega$ is closed. $d_{p}$ extends to local sections of $\wedge^{2}_{p}Z\otimes p^{*}\mathcal{O}(2)$ as we will now explain. First note that any section of $\wedge^{2}_{p}Z\otimes p^{*}\mathcal{O}(2)$ may be written locally as a product of the form $\omega\otimes p^{*}\mu$. We then define
@@
$\Omega=\frac{\Omega_{-}}{\hbar^{2}}+\frac{i\Omega_{I}}{\hbar}+\Omega_{+}$ (2.9)
@@
Given such an $\Omega$, we have for any vector field $V$ in the distribution $L$, $\mathcal{L}_{V}\Omega=0$, using Cartan’s formula, so $\Omega$ descends to a $2$-form on $Z$ and defines a relative $2$-form by restriction to $\operatorname{ker}dp$. Take the tensor product with the section of $p^{*}\mathcal{O}(2)$ corresponding to the polynomial $u^{1}u^{1}$. The resulting relative $\mathcal{O}(2)$-valued $2$-form is $d_{p}$-closed from the definition of $d_{p}$ by (2.8). Conversely, given a closed $\mathcal{O}(2)$-valued relative $2$-form trivialise the bundle $p^{*}\mathcal{O}(2)$ by the section corresponding to $u^{1}u^{1}$ to obtain a section $\omega$ of $\wedge^{2}_{p}Z$. Vector fields on $Z$ correspond to (equivalence classes of) vector fields $U$ on $X\times\mathbb{CP}^{1}$ commuting with $L$ in the sense $U,L]\subseteq L$. The subbundle $\operatorname{ker}dp$ corresponds to those $U$ with no $\partial_{\hbar}$ terms. There is therefore a unique $2$-form $\Omega$ on $X\times\mathbb{CP}^{1}$ annihilating tangents to the second factor, descending to a $2$-form on $Z$ which restricts to $\omega$. Since the section corresponding to $u^{1}u^{1}$ has a zero of order two at $\hbar=0$, in order the tensor product of this section and $\omega$ is holomorphic, $\Omega$ necessarily takes the form (2.9). $\omega$ is relatively closed so the exterior derivative of $\Omega$ vanishes on restriction to tangents to the first factor, hence $\Omega_{-},\Omega_{I},\Omega_{+}$ must be all closed. These constructions are clearly inverse to one another. ∎
A closed $\mathcal{O}(2)$-valued $2$-form is of maximal rank if it induces a non-degenerate form on each fibre of the fibration $Z\to\mathbb{CP}^{1}$. In terms of $\Omega$ this is equivalent to $\mbox{dim}\ker\Omega=2n+1$ or alternatively $\ker\Omega_{-}=\text{span}\{V_{a}\}_{a=1}^{2n},\ker\Omega_{+}=\text{span}\{U_{a}\}_{a=1}^{2n}$.
@@
Given a complex hyper-Kähler structure on $X$ we may construct a twistor distribution by choosing a trivialisation $\{U_{a}\}_{a=1}^{2n}$ for the $i$ eigenspace of $I$, defining $V_{a}\coloneqq J(U_{a})$ and defining the twistor distribution as per (2.2). It is integrable for all $\hbar$ due to the integrability of the complex structures (which follows from the complex structures being parallel for the Levi-Civita connection). We now define
$\Omega_{-}$ $\coloneqq\frac{1}{2}\big{(}g(J\cdot,\cdot)-ig(K\cdot,\cdot)\big{)}$ (2.10)
$\Omega_{I}$ $\coloneqq g(I\cdot,\cdot)$ (2.11)
@@
With these definitions $\Omega$ as in (2.9) defines a closed relative $\mathcal{O}(2)$-valued $2$-form on the twistor space of maximal rank.
Conversely, given a closed relative $\mathcal{O}(2)$-valued $2$-form on the twistor space of maximal rank. We define
@@
That this is non-degenerate follows from the fact $\Omega_{I}$ is non-degenerate. To see this, one should consider the equations
$\Omega_{-}(U_{a},\cdot)+\hskip 1.13809pti\Omega_{I}(V_{a},\cdot)$ $=0$ (2.14)
@@
that arise from $\Omega$ annihilating $L$ for each $\hbar$ and then note that this shows $\Omega_{I}$ defines an isomorphism from $TX$ to $\wedge^{1}X$ since the $\Omega_{-}(U_{a},\cdot)$ and $\Omega_{+}(V_{a},\cdot)$ give $4n$-linearly independent $1$-forms a=1,,2na=1,...,2n. Next, (2.14, 2.15) together imply
$-(\Omega_{-}+\Omega_{+})(J\cdot,\cdot)=g=-(\Omega_{-}-\Omega_{+})(K\cdot,\cdot)$ (2.16)
from which hermicity with respect to $I,J,K$ follows immediately. Lastly, the integrability of the complex structures together with closure of the $2$-forms $\Omega_{-},\Omega_{I},\Omega_{+}$ implies that the complex structures are parallel. ∎
@@ 3. Joyce structures
A geometric axiomatisation of Joyce structures is provided in [12]. Let $(M,\omega)$ be a complex manifold of complex dimension $2n$ with holomorphic symplectic form $\omega$.  [^1] To avoid confusion, in [12], Joyce structures compatible with a non-degenerate symplectic form as we will assume here are called strong Joyce structures. Let $TM$ be the holomorphic tangent bundle with equipped with the natural projection $\pi:TM\to M$. A lattice $\mathcal{G}_{p}$ at $p$ is a discrete subgroup of $T_{p}M$ such that complex multiplication $\mathcal{G}_{p}\otimes_{\mathbb{Z}}\mathbb{C}\to TM$ is an isomorphism. A holomorphic bundle of lattices $\mathcal{G}\hookrightarrow TM$ defines a flat linear connection $\nabla$ on $TM$ by distinguishing the $2n$-dimensional vector space of parallel sections. Specifically, the vector space of parallel sections is generated by those sections taking values in the lattice.
@@
Let $M$ be a complex manifold of dimension $2n$. A period structure on $M$ is a holomorphic bundle of lattices $\mathcal{G}\to TM$ together with a holomorphic symplectic form $\omega$ on $M$ that takes rational values on the lattice and a holomorphic vector field $E_{0}$ such that $\nabla E_{0}=\operatorname{Id}_{TM}$ where $\nabla$ is the flat connection associated to the lattice.
@@
Let $M$ be a complex manifold of dimension $2n$. Let $\pi:TM\to M$ be the natural projection. A (strong) Joyce structure is a period structure with symplectic form on $M$ together with a complex hyper-Kähler structure $(g,I,J,K)$ on the total space $TM$ satisfying the following:
@@
    $\pi^{*}\omega=\Omega_{-}$, where $\Omega_{-}\coloneqq(g(J\cdot,\cdot)-ig(K\cdot,\cdot))/2$
@@ 4. Quadratic differentials and the moduli space $M={\rm Quad}(\bar{m})$
A meromorphic quadratic differential on $\mathbb{CP}^{1}$ is a meromorphic section of $K_{\mathbb{CP}^{1}}^{2}$. With respect to the affine coordinate $x$ for $\mathbb{CP}^{1}$ we may write any such quadratic differential like $\phi=\mathcal{Q}(x)dx^{2}$ for a rational function $\mathcal{Q}(x)$ on $\mathbb{C}$.
@@
Given the conditions above we are assured that $M$ is non-empty of dimension greater than or equal to two. The condition that there is a pole of order greater than or equal to five is mainly for convenience in choosing local coordinates. Indeed, later we will handle separately the case with four simple poles, along the same lines as the restricted class above. Note that the moduli spaces considered here include the cases of $\bar{m}=\{2n+5\}$ for $n\geq 1$ which were considered in [15].
@@
$Q_{0}(x)$ $=x^{2m_{\infty}-5}+\sum_{i=0}^{2m_{\infty}-7}a^{(\infty)}_{i}x^{i}+\sum_{\alpha=1}^{N}\sum_{i=1}^{2m_{\alpha}-1}\frac{a_{i}^{(\alpha)}}{(x-w_{\alpha})^{i}}.$ (4.1)
@@
and thinking of $\phi$ as a holomorphic section of $K_{\mathbb{CP}^{1}}(D)\otimes K_{\mathbb{CP}^{1}}(D)$. More practically, the affine piece is realised as a subspace of $\mathbb{C}^{2}$ by the equation $y_{0}^{2}=Q_{0}(x)$. The covering map $\Pi:\Sigma_{0}\to\mathbb{CP}^{1}$ is branched precisely at the poles and zeroes of $\phi$. There is a tautological $1$-form $\psi$, a meromorphic section of the canonical bundle $K_{\Sigma}$ defined (up to a choice of sign) by the condition $\Pi^{*}\phi=\psi\otimes\psi$. It has local representation $\psi=y_{0}\,dx$ in the charts for $\Sigma$ induced by the chart for $\mathbb{CP}^{1}$ in which we may write $\phi=Q_{0}(x)dx^{2}$. $\psi$ has poles at the pre-images under $\Pi$ of poles of $\phi$ of order greater than one. Precisely: $\psi$ has a pole of order $2m_{\alpha}-2$ at the branch point $w_{\alpha}$ (note that if $\phi$ has a simple pole at $w_{\alpha}$ then $\psi$ will be holomorphic at the corresponding branch point). The $1$-form has no residues at its poles since pulling back by the covering involution $y\mapsto-y$ sends $\psi\mapsto-\psi$. An important consequence of the vanishing residues is that $\psi$ defines an element $\psi]\in H^{1}(\Sigma_{0},\mathbb{C})$.
@@
Here $U(y_{0})$ is defined as the derivative of the function $y_{0}$ (thought of a function of $x$ and the coordinates on $M$) in the direction $U$. For brevity we will adopt the notation:
@@
We will see that $\mu_{0}$ is an isomorphism use it to equip $M$ with a period structure.
@@
where $\alpha,\beta$ are smooth representatives. Since $\{\gamma_{a}\}_{a=1}^{2n}$ is a canonical basis $\langle\nu_{a}|\nu_{b}\rangle_{0}=\omega_{ab}$ where $\omega_{ab}$ is the standard symplectic matrix of rank $2n$.
@@
$\omega\coloneqq\mu_{0}^{*}\langle\cdot|\cdot\rangle_{0}$ (4.11)
@@
$\omega(U,V)=2\pi i\hskip-5.69054pt\sum_{x\in\mathcal{P}_{0}(p)}\,\underset{x}{\operatorname{Res}}(d^{-1}U(\psi)V(\psi))$ (4.12)
@@
$\mu_{0}(U)=[U(\psi)]=[U(\psi)-d(\rho d^{-1}U(\psi))]$ (4.13)
Since the wedge product of holomorphic $1$-forms vanishes
@@
$=\oint_{\partial S}d^{-1}U(\psi)V(\psi)$ (4.14)
where we have used Stokes’ theorem and that $\rho=1$ on $\partial S$ in the last step. Here the integration over $\partial S$ is taken with respect to the usual orientation. By Cauchy’s integral formula
@@
$+\omega([U,V],W)-\omega([U,W],V)-\omega([V,W],U),$ (4.16)
@@
$\bigg{\{}\frac{\partial}{\partial a_{i}^{(\alpha)}}\bigg{\}}_{\alpha=1,...,N}^{i=1,...,2m_{\alpha}}\cup\bigg{\{}\frac{\partial}{\partial a_{i}^{(\infty)}}\bigg{\}}^{i=1,...,2m_{\infty}-7}$ (4.17)
@@
The functions $\{z_{a}\}_{a=1}^{2n}$ give local coordinates on $M$.

### v2

@@ 1. Introduction
Joyce structures were proposed by Bridgeland in [9] as a way of geometrising the study of Donaldson-Thomas invariants. They are expected to exist on the complex manifold $M$ of stability conditions of a $CY_{3}$ triangulated category satisfying certain assumptions. Some properties of these structures were earlier derived in [21] and [22]. The argument for their existence involves solving Riemann-Hilbert problems with infinite dimensional gauge group, defined by the variation of the Donaldson-Thomas invariants and as far as the author understands, solutions to these problems are not yet fully understood. A geometric axiomatisation of Joyce structures without reference to their origin in Donaldson-Thomas theory is provided in [13] in terms of a complex hyper-Kähler metric on the total space $TM$ satisfying certain symmetries compatible with an integral structure on $M$.
@@
The symplectic geometry and Hamiltonian structure of isomonodromic deformations have long been studied [4] and are still of active interest [3]. The infinitesimal generators may be obtained as the kernel of a fundamental 2-form, and formulae in a general setting are provided in [31]. In this work, within the setting of second order linear ODEs in one dependent variable, we take a (at least superficially) different approach and construct a fundamental $2$-form in terms of the intersection pairings of a family of algebraic curves.
We will calculate isomonodromic deformations of a large class of ODEs in the complex variable $x$ of the form
$\frac{d^{2}Y}{dx^{2}}=Q(x)Y.$ (1.1)
Specifically, we consider those such that the potential takes the form
$Q(x)=\frac{Q_{0}(x)}{\hbar^{2}}+\frac{Q_{1}(x)}{\hbar}+Q_{2}(x),$ (1.2)
with $Q_{0}(x),Q_{1}(x),Q_{2}(x)$ rational functions depending on a point $\xi\in X$, where $X$ is a complex manifold, and $\hbar\in\mathbb{C}^{*}$ is the spectral parameter.
@@ 2. Complex hyper-Kähler metrics and their twistor theory
Let $X$ be a complex manifold of dimension $4n$. A complex hyper-Kähler structure is a holomorphically varying, non-degenerate bilinear form on $TX$, the holomorphic tangent bundle of $X$, together with holomorphic endomorphisms $I,J,K$ of $TX$ satisfying
@@
$I^{*}g$ $=J^{*}g=K^{*}g=g,$
$\nabla I$ $=\nabla J=\nabla K=0,$
@@
Crucial to constructing our metrics will be families of rank-$2n$ distributions (subbundles of $TX$) depending on a spectral parameter $\hbar=u^{1}/u^{0}\in\mathbb{C}^{*}$ where $u^{0}:u^{1}]\in\mathbb{CP}^{1}$.
@@
$L(\hbar)=\operatorname{span}\Big\{L_{a}\coloneqq U_{a}+\frac{V_{a}}{\hbar}\Big\}_{a=1}^{2n},$ (2.2)
@@
$g=\sum_{a=1}^{2n}\sum_{b=1}^{2n}\omega_{ab}(U^{a}\otimes V^{b}+V^{b}\otimes U^{a}),$ (2.4)
@@
$L(\hbar)$ Frobenius integrable for all $\hbar$ is equivalent (see Appendix A) to integrability of the $\pm i$-eigenspaces of the complex structures $I,J,K$. We may think of $L$ as a subbundle of $T(X\times\mathbb{CP}^{1})$. Then, integrability implies we may construct the space of leaves of this distribution.
@@
$\wedge^{k}_{p}T^{*}Z\coloneqq\wedge^{k}({\operatorname{ker}dp})^{*}$ (2.6)
denote the bundle of $k$-forms relative to $p$. A relative (to $p$) $\mathcal{O}(2)$-valued 2-form is a section of the bundle $\wedge^{2}_{p}T^{*}Z\otimes p^{*}\mathcal{O}(2)$.
We have the relative exterior derivative $d_{p}:\Gamma(\wedge^{2}_{p}T^{*}Z)\to\Gamma(\wedge^{3}_{p}T^{*}Z)$ given by
$d_{p}\omega=d\tilde{\omega}|_{\ker dp},$ (2.7)
where $\tilde{\omega}$ is any $2$-form that restricts to $\omega$ on $\operatorname{ker}dp$. If $d_{p}\omega=0$ we say $\omega$ is closed. $d_{p}$ extends to local sections of $\wedge^{2}_{p}T^{*}Z\otimes p^{*}\mathcal{O}(2)$ as we will now explain. First note that any section of $\wedge^{2}_{p}T^{*}Z\otimes p^{*}\mathcal{O}(2)$ may be written locally as a product of the form $\omega\otimes p^{*}\mu$. We then define
@@
$\Omega=\frac{\Omega_{-}}{\hbar^{2}}+\frac{i\Omega_{I}}{\hbar}+\Omega_{+},$ (2.9)
@@
Let $q:X\times\mathbb{CP}^{1}\to Z$ be the quotient map. Suppose we have such an $\Omega$. Using Cartan’s formula, we have for a vector field $V$ in the distribution $L$, and $X,Y$ that push forward to vectors in $\ker dp$, $(\mathcal{L}_{V}\Omega)(X,Y)=0$, so $\Omega$ defines a relative $2$-form $\omega$ on $Z$. Take the tensor product with the section of $p^{*}\mathcal{O}(2)$ corresponding to the polynomial $u^{1}u^{1}$ to obtain relative $\mathcal{O}(2)$-valued $2$-form. Now, for any $2$-form $\tilde{\omega}$ on $Z$ restricting to $\omega$ on $\ker dp$, it is evident $q^{*}\tilde{\omega}$ agrees with $\Omega$ up to a $2$-form vanishing when restricted to $X$. Since $d\Omega$ vanishes when restricted to $X$, $d\tilde{\omega}|_{\ker dp}=0$, and hence the relative $\mathcal{O}(2)$-valued $2$-form is $d_{p}$-closed.
Conversely, given a closed $\mathcal{O}(2)$-valued relative $2$-form trivialise the bundle $p^{*}\mathcal{O}(2)$ by the section corresponding to $u^{1}u^{1}$ to obtain a section $\omega$ of $\wedge^{2}_{p}T^{*}Z$. We may define a $2$-form $\Omega$ on $X\times\mathbb{CP}^{1}$ as $\Omega(X,Y)=\omega(dq(X),dq(Y))$ for $X,Y$ tangent to the first factor and insisting $\Omega$ annihilate tangents to the second. Since the section corresponding to $u^{1}u^{1}$ has a zero of order two at $\hbar=0$, in order for the tensor product of this section and $\omega$ to be holomorphic, $\Omega$ necessarily takes the form (2.9). $\omega$ is relatively closed so the exterior derivative of $\Omega$ vanishes on restriction to tangents to the first factor, hence $\Omega_{-},\Omega_{I},\Omega_{+}$ must be all closed. These constructions are clearly inverse to one another. ∎
A closed $\mathcal{O}(2)$-valued $2$-form is of maximal rank if it induces a non-degenerate form on each fibre of the fibration $Z\to\mathbb{CP}^{1}$. In terms of $\Omega$ this is equivalent to $\mbox{dim}\ker\Omega=2n+1$.
@@
Given a complex hyper-Kähler structure on $X$, we may construct a twistor distribution by choosing a trivialisation $\{U_{a}\}_{a=1}^{2n}$ for the $i$-eigenspace of $I$, defining $V_{a}\coloneqq J(U_{a})$ and defining the twistor distribution as per (2.2). The complex structures (2.3) associated with the twistor distribution agree with the ones coming from the hyper-Kähler structure. $L(\hbar)$ is integrable for all $\hbar$ due to the integrability of the complex structures (which follows from the complex structures being parallel for the Levi-Civita connection). We now define
$\Omega_{-}$ $\coloneqq\frac{1}{2}\big(g(J\cdot,\cdot)-ig(K\cdot,\cdot)\big),$ (2.10)
$\Omega_{I}$ $\coloneqq g(I\cdot,\cdot),$ (2.11)
@@
With these definitions, $\Omega$ as in (2.9) defines a closed relative $\mathcal{O}(2)$-valued $2$-form on the twistor space of maximal rank.
Conversely, given a closed relative $\mathcal{O}(2)$-valued $2$-form on the twistor space of maximal rank, recall we have the quaternionic structure (2.3) associated to the twistor distribution. We define the symmetric bilinear form
@@
We consider the equations
$\Omega_{-}(U_{a},\cdot)+\hskip 1.13809pti\Omega_{I}(V_{a},\cdot)$ $=0,$ (2.14)
@@
that arise from $\Omega$ annihilating $L(\hbar)$ for each $\hbar$. Since $\mbox{dim}\ker\Omega=2n+1$, we have $\ker\Omega_{-}=\text{span}\{V_{a}\}_{a=1}^{2n}$ and $\ker\Omega_{+}=\text{span}\{U_{a}\}_{a=1}^{2n}$. This implies $\Omega_{-}(U_{a},\cdot)$ and $\Omega_{+}(V_{a},\cdot)$ give $4n$-linearly independent $1$-forms a=1,,2na=1,...,2n. Thus (2.14, 2.15) imply $\Omega_{I}$ defines an isomorphism from $TX$ to $T^{*}X$ and hence that $g$ defines a non-degenerate bilinear form. Next, (2.14, 2.15) together imply
$-(\Omega_{-}+\Omega_{+})(J\cdot,\cdot)=g=-i(\Omega_{-}-\Omega_{+})(K\cdot,\cdot),$ (2.16)
from which Hermiticity with respect to $I,J,K$ follows immediately. Lastly, the integrability of the complex structures together with closure of the $2$-forms $\Omega_{-},\Omega_{I},\Omega_{+}$ implies that the complex structures are parallel. ∎
@@ 3. Joyce structures
A geometric axiomatisation of Joyce structures is provided in [13]. Let $(M,\omega)$ be a complex manifold of complex dimension $2n$ with holomorphic symplectic [^1] To avoid confusion, in [13], Joyce structures compatible with a non-degenerate symplectic form as we will assume here are called strong Joyce structures. form $\omega$. Let $TM$ be the holomorphic tangent bundle equipped with the natural projection $\pi:TM\to M$. A lattice $\mathcal{G}_{p}$ at $p$ is a discrete subgroup of $T_{p}M$ such that complex multiplication $\mathcal{G}_{p}\otimes_{\mathbb{Z}}\mathbb{C}\to TM$ is an isomorphism. A holomorphic bundle of lattices $\mathcal{G}\hookrightarrow TM$ defines a flat linear connection $\nabla$ on $TM$ by distinguishing the $2n$-dimensional vector space of parallel sections. Specifically, the vector space of parallel sections is generated by those sections taking values in the lattice.
@@
Let $M$ be a complex manifold of dimension $2n$. A period structure on $M$ is a holomorphic bundle of lattices $\mathcal{G}\to TM$ together with a holomorphic symplectic form $\omega$ on $M$ that takes rational values on the lattice, and a holomorphic vector field $E_{0}$ such that $\nabla E_{0}=\operatorname{Id}_{TM}$ where $\nabla$ is the flat connection associated to the lattice.
@@
Let $M$ be a complex manifold of dimension $2n$. Let $\pi:TM\to M$ be the natural projection. A (strong) Joyce structure is a period structure with symplectic form $\omega$ on $M$ together with a complex hyper-Kähler structure $(g,I,J,K)$ on the total space $TM$ satisfying the following:
@@
    $\pi^{*}\omega=\Omega_{-}$, where $\Omega_{-}\coloneqq(g(J\cdot,\cdot)-ig(K\cdot,\cdot))/2$.
@@ 4. Quadratic differentials and the moduli space $M={\rm Quad}(\bar{m})$
A meromorphic quadratic differential on $\mathbb{CP}^{1}$ is a meromorphic section of $K_{\mathbb{CP}^{1}}^{2}$. With respect to the affine coordinate $x$ for $\mathbb{CP}^{1}$ we may write any such quadratic differential as $\phi=\mathcal{Q}(x)dx^{2}$ for a rational function $\mathcal{Q}(x)$ on $\mathbb{C}$.
@@
Given the conditions above, we are assured that $M$ is non-empty of dimension greater than or equal to two. The condition that there is a pole of order greater than or equal to five is mainly for convenience in choosing local coordinates. Indeed, later we will handle separately the case with four simple poles, along the same lines as the restricted class above. Note that the moduli spaces considered here include the cases of $\bar{m}=\{2n+5\}$ for $n\geq 1$ which were considered in [16].
@@
$Q_{0}(x)$ $=x^{2m_{\infty}-5}+\sum_{i=0}^{2m_{\infty}-7}a^{(\infty)}_{i}x^{i}+\sum_{\alpha=1}^{N}\sum_{i=1}^{2m_{\alpha}-1}\frac{a_{i}^{(\alpha)}}{(x-w_{\alpha})^{i}},$ (4.1)
@@
and thinking of $\phi$ as a holomorphic section of $K_{\mathbb{CP}^{1}}(D)\otimes K_{\mathbb{CP}^{1}}(D)$. More practically, the affine piece is realised as a subspace of $\mathbb{C}^{2}$ by the equation $y_{0}^{2}=Q_{0}(x)$ away from the poles. The covering map $\Pi:\Sigma_{0}\to\mathbb{CP}^{1}$ is branched precisely at the poles and zeroes of $\phi$. There is a tautological $1$-form $\psi$, a meromorphic section of the canonical bundle $K_{\Sigma}$ defined (up to a choice of sign) by the condition $\Pi^{*}\phi=\psi\otimes\psi$. It has local representation $\psi=y_{0}\,dx$ in the charts for $\Sigma$ induced by the chart for $\mathbb{CP}^{1}$ in which we may write $\phi=Q_{0}(x)dx^{2}$. $\psi$ has poles at the pre-images under $\Pi$ of poles of $\phi$ of order greater than one. Precisely: $\psi$ has a pole of order $2m_{\alpha}-2$ at the branch point $w_{\alpha}$ (note that if $\phi$ has a simple pole at $w_{\alpha}$ then $\psi$ will be holomorphic at the corresponding branch point). The $1$-form has no residues at its poles since pulling back by the covering involution $y\mapsto-y$ sends $\psi\mapsto-\psi$. An important consequence of the vanishing residues is that $\psi$ defines an element $\psi]\in H^{1}(\Sigma_{0},\mathbb{C})$.
@@
Here $U(y_{0})$ is defined as the derivative of the function $y_{0}$ (thought of as a function of $x$ and the coordinates on $M$) in the direction $U$. For brevity we will adopt the notation:
@@
We will see that $\mu_{0}$ is an isomorphism and use it to equip $M$ with a period structure.
@@
where $\alpha,\beta$ are smooth representatives. Since $\{\gamma_{a}\}_{a=1}^{2n}$ is a canonical basis, $\langle\nu_{a}|\nu_{b}\rangle_{0}=\omega_{ab}$ where $\omega_{ab}$ is the standard symplectic matrix of rank $2n$.
@@
$\omega\coloneqq\mu_{0}^{*}\langle\cdot|\cdot\rangle_{0},$ (4.11)
@@
$\omega(U,V)=2\pi i\hskip-5.69054pt\sum_{x\in\mathcal{P}_{0}(p)}\,\underset{x}{\operatorname{Res}}(d^{-1}U(\psi)V(\psi)),$ (4.12)
@@
$\mu_{0}(U)=[U(\psi)]=[U(\psi)-d(\rho d^{-1}U(\psi))].$ (4.13)
Since the wedge product of holomorphic $1$-forms vanishes,
@@
$=\oint_{\partial S}d^{-1}U(\psi)V(\psi),$ (4.14)
where we have used Stokes’ theorem and that $\rho=1$ on $\partial S$ in the last step. Here the integration over $\partial S$ is taken with respect to the usual orientation. By Cauchy’s integral formula,
@@
$+\omega([U,V],W)-\omega([U,W],V)-\omega([W,V],U),$ (4.16)
@@
$\bigg\{\frac{\partial}{\partial a_{i}^{(\alpha)}}\bigg\}_{\alpha=1,...,N}^{i=1,...,2m_{\alpha}}\cup\bigg\{\frac{\partial}{\partial a_{i}^{(\infty)}}\bigg\}^{i=0,...,2m_{\infty}-7}$ (4.17)
@@
The functions $\{z^{a}\}_{a=1}^{2n}$ give local coordinates on $M$.